# The inducibility of Turán graphs

### v1

@@ 5 Exact results
In this section, we show that extremal graphs for $I(F,n)$ are Turán graphs whenever $F$ is an almost balanced complete multipartite graph, thus proving Theorem 1.1. First, we prove a technical lemma.
@@
Let $s,t\geq 2$ be integers. Suppose $x,y,N$ are real numbers with $y>x+1$ satisfying |xN|=o(N)|x-N|=o(N) and |yN|=o(N)|y-N|=o(N) as $N\to\infty$. Then
$\tbinom{x+1}{s}+\tbinom{y-1}{s}-\tbinom{x}{s}-\tbinom{y}{s}=-\tfrac{1}{(s-2)!}(y-x-1+o(1))N^{s-2},$
@@
$\tbinom{x+1}{t}\tbinom{y-1}{s}+\tbinom{x+1}{s}\tbinom{y-1}{t}-\tbinom{x}{t}\tbinom{y}{s}-\tbinom{x}{s}\tbinom{y}{t}=\tfrac{t+s-(t-s)^{2}}{s!t!}\big(y-x-1+o(1)\big)N^{t+s-2}.$
@@
By the Mean Value Theorem, there exists $z_{0}$ between $x$ and y1y-1 (hence $|z_{0}-N|=o(N)$) such that
$\tbinom{x+1}{s}+\tbinom{y-1}{s}-\tbinom{x}{s}-\tbinom{y}{s}$ $=\tbinom{x}{s-1}-\tbinom{y-1}{s-1}=(x-y+1)\bigl(\tbinom{z}{s-1}\bigr)^{\prime}\big|_{z=z_{0}}$
$=\tfrac{1}{(s-2)!}(x-y+1)\bigl(z_{0}^{s-2}+o(z_{0}^{s-2})\bigr)=-\tfrac{1}{(s-2)!}\bigl(y-x-1+o(1)\bigr)N^{s-2}.$
For the second formula, let
A direct computation gives
Applying the Mean Value Theorem to the two differences, we find $z_{1},z_{2}$ between $x$ and y1y-1 with $|z_{i}-N|=o(N)$ such that
$y(y-s)_{t-s}-(x+1)(x+1-s)_{t-s}$ $=(y-x-1)\bigl(z(z-s)_{t-s}\bigr)^{\prime}\big|_{z=z_{1}}$
$=\bigl(y-x-1+o(1)\bigr)(t-s+1)N^{t-s},$
$\tfrac{(y-s)_{t-s}}{y}-\tfrac{(x+1-s)_{t-s}}{x+1}$ $=(y-x-1)\Bigl(\tfrac{(z-s)_{t-s}}{z}\Bigr)^{\prime}\big|_{z=z_{2}}$
$=\bigl(y-x-1+o(1)\bigr)(t-s-1)N^{t-s-2}.$
Since $x,y=N+o(N)$, we have $t(x+1)y=tN^{2}+o(N^{2})$. Substituting these asymptotics into $g(x,y)$ yields
$g(x,y)=\bigl(y-x-1+o(1)\bigr)\bigl(t+s-(t-s)^{2}\bigr)N^{t-s}.$
Finally, using $(x)_{s-1}=N^{s-1}+o(N^{s-1})$ and $(y-1)_{s-1}=N^{s-1}+o(N^{s-1})$,
$\tbinom{x+1}{t}\tbinom{y-1}{s}+\tbinom{x+1}{s}\tbinom{y-1}{t}-\tbinom{x}{t}\tbinom{y}{s}-\tbinom{x}{s}\tbinom{y}{t}$
$\quad=\frac{(x)_{s-1}(y-1)_{s-1}}{t!\,s!}\,g(x,y)=\frac{t+s-(t-s)^{2}}{s!\,t!}\,\bigl(y-x-1+o(1)\bigr)N^{t+s-2}.$
This completes the proof. ∎
@@
Let $F$ be an almost balanced complete $r$-partite graph on $\ell$ vertices with part sizes $a_{1}\geq\cdots\geq a_{r}$. Let $n$ be a sufficiently large integer and set $m\coloneqq m_{r,\ell}$. By Theorem 3.1 and Theorem 1.5, every $n$-vertex graph $G$ attaining $I(F,n)$ is a complete $m$-partite graph whose parts have sizes $(1+o(1))N$, where $N\coloneqq n/m$.
We now determine the exact part sizes. Write the part sizes of $G$ as $n_{1}\geq n_{2}\geq\dots\geq n_{m}$, with $\sum_{i=1}^{m}n_{i}=n$ and $|n_{i}-N|=o(n)$ for all $i$. Let $y\coloneqq n_{1}$ and $x\coloneqq n_{m}$; by definition $y\geq x$.
@@
$\binom{\ell}{2}>m\sum_{k=1}^{r}\binom{a_{k}}{2}$.
@@
Since $F$ is almost balanced, we have $(a_{i}-a_{j})^{2}\leq a_{i}+a_{j}-2$ for all $1\leq i<j\leq r$. Therefore,
$r\sum_{k=1}^{r}{a_{k}}^{2}=\Bigl(\sum_{k=1}^{r}a_{k}\Bigr)^{2}+\sum_{1\leq i<j\leq r}(a_{i}-a_{j})^{2}\leq\ell^{2}+\sum_{1\leq i<j\leq r}(a_{i}+a_{j}-2)=\ell^{2}+(r-1)\ell-r(r-1).$
Consequently,
$r\sum_{k=1}^{r}\binom{a_{k}}{2}\leq\binom{\ell}{2}-\binom{r}{2}$. (37)
By Lemma 4.2 (ii), we have
$m<\frac{r\ell^{2}}{\ell^{2}-r^{2}}=\frac{r\ell^{2}}{(\ell-r)(\ell+r)}<\frac{r\ell(\ell-1)}{(\ell-r-1)(\ell+r)}=\frac{r\binom{\ell}{2}}{\binom{\ell}{2}-\binom{r}{2}}$. (38)
Combining (37) and (38) yields $\binom{\ell}{2}>m\sum_{k=1}^{r}\binom{a_{k}}{2}$. ∎
For $\{i,j\}\subseteq[r]$, define
@@
$A_{i}$ $\coloneqq\sum_{(s_{k})_{k\in[r]\setminus\{i\}}}\prod_{k\in[r]\setminus\{i\}}\tbinom{n_{s_{k}}}{a_{k}}=\left(\frac{(m-2)_{r-1}\cdot a_{i}!}{\prod_{k\in[r]}a_{k}!}+o(1)\right)N^{\ell-a_{i}},$
@@
$B_{i,j}$ $\coloneqq\sum_{(s_{k})_{k\in[r]\setminus\{i,j\}}}\prod_{k\in[r]\setminus\{i,j\}}\tbinom{n_{s_{k}}}{a_{k}}=\left(\frac{(m-2)_{r-2}\cdot a_{i}!a_{j}!}{\prod_{k\in[r]}a_{k}!}+o(1)\right)N^{\ell-a_{i}-a_{j}},$
where the tuples $(s_{k})$ run over $\big([2,m-1]\big)_{r-1}$ for $A_{i}$ and run over $\big([2,m-1]\big)_{r-2}$ for $B_{i,j}$ (the indices $1$ and $m$ correspond to the parts of sizes $y$ and $x$, which will be treated separately).
Set $\hat{I}(F,H)\coloneqq\mathrm{sym}(a_{1},\dots,a_{r})\cdot I(F,H)$ for any graph $H$. Let $G_{*}\cong K_{n_{1}-1,n_{2},\dots,n_{m}+1}$. By (4),
$\hat{I}(F,G_{*})-\hat{I}(F,G)$ $=\sum_{i\in[r]}\left(\tbinom{x+1}{a_{i}}+\tbinom{y-1}{a_{i}}-\tbinom{x}{a_{i}}-\tbinom{y}{a_{i}}\right)A_{i}$
$\quad+\sum_{\{i,j\}\subseteq[r]}\left(\tbinom{x+1}{a_{i}}\tbinom{y-1}{a_{j}}+\tbinom{x+1}{a_{j}}\tbinom{y-1}{a_{i}}-\tbinom{x}{a_{i}}\tbinom{y}{a_{j}}-\tbinom{x}{a_{j}}\tbinom{y}{a_{i}}\right)B_{i,j}.$ (39)
If $a_{i}=1$, the coefficient of $A_{i}$ vanishes. Otherwise $a_{i}\geq 2$, Lemma 5.1 gives
$\left(\tbinom{x+1}{a_{i}}+\tbinom{y-1}{a_{i}}-\tbinom{x}{a_{i}}-\tbinom{y}{a_{i}}\right)A_{i}$
$={}$ $-\frac{1}{(a_{i}-2)!}\bigl(y-x-1+o(1)\bigr)N^{a_{i}-2}\cdot\left(\frac{(m-2)_{r-1}\cdot a_{i}!}{\prod_{k\in[r]}a_{k}!}+o(1)\right)N^{\ell-a_{i}}$
$={}$ $-\frac{\bigl(y-x-1+o(1)\bigr)(m-2)_{r-1}}{\prod_{k\in[r]}a_{k}!}\cdot a_{i}(a_{i}-1)\,N^{\ell-2}.$ (40)
(Note that when $a_{1}=1$, (5) holds trivially because both sides are zero.) Similarly, Lemma 5.1 yields
$\left(\tbinom{x+1}{a_{i}}\tbinom{y-1}{a_{j}}+\tbinom{x+1}{a_{j}}\tbinom{y-1}{a_{i}}-\tbinom{x}{a_{i}}\tbinom{y}{a_{j}}-\tbinom{x}{a_{j}}\tbinom{y}{a_{i}}\right)B_{i,j}$
$={}$ $\frac{a_{i}+a_{j}-(a_{i}-a_{j})^{2}}{a_{i}!a_{j}!}\bigl(y-x-1+o(1)\bigr)N^{a_{i}+a_{j}-2}\cdot\left(\frac{(m-2)_{r-2}\cdot a_{i}!a_{j}!}{\prod_{k\in[r]}a_{k}!}+o(1)\right)N^{\ell-a_{i}-a_{j}}$
$={}$ $\frac{\bigl(y-x-1+o(1)\bigr)(m-2)_{r-2}}{\prod_{k\in[r]}a_{k}!}\cdot\bigl(a_{i}+a_{j}-(a_{i}-a_{j})^{2}\bigr)\,N^{\ell-2}.$ (41)
Substituting (5) and (5) into (5) gives
$\hat{I}(F,G_{*})-\hat{I}(F,G)={}$ $\frac{\bigl(y-x-1+o(1)\bigr)(m-2)_{r-2}}{\prod_{k\in[r]}a_{k}!}N^{\ell-2}$
$\times\Biggl(-(m-r)\sum_{i\in[r]}a_{i}(a_{i}-1)+\sum_{\{i,j\}\subseteq[r]}\bigl(a_{i}+a_{j}-(a_{i}-a_{j})^{2}\bigr)\Biggr).$
Straightforward calculations show that
$-(m-r)\sum_{i\in[r]}a_{i}(a_{i}-1)+\sum_{\{i,j\}\subseteq[r]}\big(a_{i}+a_{j}-(a_{i}-a_{j})^{2}\big)=2\Bigl(\tbinom{\ell}{2}-m\sum_{i\in[r]}\tbinom{a_{i}}{2}\Bigr)$.
Therefore,
$\hat{I}(F,G_{*})-\hat{I}(F,G)=\frac{2\bigl(y-x-1+o(1)\bigr)(m-2)_{r-2}}{\prod_{k\in[r]}a_{k}!}\cdot N^{\ell-2}\Bigl(\tbinom{\ell}{2}-m\sum_{i\in[r]}\tbinom{a_{i}}{2}\Bigr)$. (42)
By Claim 5.2, the factor $\binom{\ell}{2}-m\sum_{i}\binom{a_{i}}{2}$ is positive. If $y>x+1$, then for sufficiently large $n$ the term yx1+o(1)y-x-1+o(1) is also positive, implying $\hat{I}(F,G_{*})>\hat{I}(F,G)$. This shows that $I(F,G_{*})>I(F,G)$ as well, contradicting the maximality of $G$. Hence we must have $y\leq x+1$, which forces all part sizes of $G$ to differ by at most $1$, i.e., $G$ is a Turán graph $T_{m}(n)$. This completes the proof of Theorem 1.1. ∎
@@
Let $r,t\geq 2$ be integers. For every integer $n$, there exists an integer $k$ such that the Turán graph $T_{k}(n)$ is an extremal graph for $I(K_{r}(t),n)$.
@@
The case $n<rt$ is trivial. For $n\geq rt$, Lemma 2.2 guarantees that some complete multipartite graph $G$ is extremal for $I(K_{r}(t),n)$. Let $G$ have $k$ parts with sizes $a_{1}\geq\dots\geq a_{k}$. Since $G$ must accommodate an induced copy of $K_{r}(t)$, we necessarily have $k\geq r$ and $a_{k}\geq t$.
Set $F\coloneqq K_{r}(t)$. Define $G^{\prime}$ as the complete $k$-partite graph with part sizes $a_{1}-1,a_{2},\dots,a_{k-1},a_{k}+1$, and $G^{\prime\prime}$ as the complete (k1)(k-1)-partite graph with part sizes $a_{1}+a_{k},a_{2},\dots,a_{k-1}$.
@@
If $a_{1}-a_{k}>1$, then $I(F,G)<\max\{I(F,G^{\prime}),I(F,G^{\prime\prime})\}$.

### v2

@@ 5 Exact results and uniqueness of the extremal construction
In this section, we show that, for large $n$, the extremal graph for $I(F,n)$ is the Turán graph if $F$ is an almost balanced complete multipartite graph, thus completing the proof of Theorem 1.1. We begin with the following two lemmas.
@@
Let $t\geq s\geq 1$ be integers. Suppose $x,y,N$ are real numbers satisfying $y>x+1$ and $\max\{|x-N|,~|y-N|\}=o(N)$ as $N\to\infty$. Then
$\tbinom{x+1}{s}+\tbinom{y-1}{s}-\tbinom{x}{s}-\tbinom{y}{s}$ $=-\tfrac{s(s-1)}{s!}(y-x-1)N^{s-2}+o(N^{s-2});$ (32)
$\tbinom{x+1}{t}\tbinom{y-1}{s}+\tbinom{x+1}{s}\tbinom{y-1}{t}-\tbinom{x}{t}\tbinom{y}{s}-\tbinom{x}{s}\tbinom{y}{t}$ $=\tfrac{t+s-(t-s)^{2}}{s!t!}\big(y-x-1\big)N^{t+s-2}+o(N^{t+s-2}).$ (33)
Proof.
We start with the first equality. The case $s=1$ is trivially true, so we may assume that $s\geq 2$. Let $\phi\colon[s-1,\infty)\to\mathbb{R}$ be the function defined by $\phi(z)=\binom{z}{s-1}$ for all $z\in[s-1,\infty)$. By the Mean Value Theorem, there exists $z_{0}\in[x,y-1]$ (hence $|z_{0}-N|=o(N)$) such that
$\tbinom{x+1}{s}+\tbinom{y-1}{s}-\tbinom{x}{s}-\tbinom{y}{s}=\tbinom{x}{s-1}-\tbinom{y-1}{s-1}$ $=(x-y+1)\phi^{\prime}(z_{0})$
$=(x-y+1)\left(\tfrac{z_{0}^{s-2}}{(s-2)!}+o\left(z_{0}^{s-2}\right)\right)=-\tfrac{y-x-1}{(s-2)!}N^{s-2}+o(N^{s-2}),$
as desired.
We now consider the second equality. Let
Claim 5.2.
We have $g(x,y)=\left(t+s-(t-s)^{2}\right)(y-x-1)N^{t-s}+o(N^{t-s})$.
Proof.
The case $t-s\in\{0,1\}$ follows directly from the following fact (which itself follows from straightforward calculations):
$g(x,y)=\begin{cases}2s(y-x-1),&\quad\text{if}\quad t-s=0,\\ s(y-x-1)(x+y-2s),&\quad\text{if}\quad t-s=1.\end{cases}$
So we may assume that $t-s\geq 2$. Simplifying $g(x,y)$, we obtain
Let $\psi_{1},\psi_{2}\colon[t-1,\infty)\to\mathbb{R}$ be functions defined by
$\psi_{1}(z)=z(z-s)_{t-s}\quad\text{and}\quad\psi_{2}(z)=\frac{(z-s)_{t-s}}{z}\quad\text{for all}\quad z\in[t-1,\infty).$
By the Mean Value Theorem, there exist $z_{1},z_{2}\in[x,y-1]$ such that
$y(y-s)_{t-s}-(x+1)(x+1-s)_{t-s}$ $=(y-x-1)\psi_{1}^{\prime}(z_{1})$
$=(y-x-1)\left((t-s+1)z_{1}^{t-s}+o(z^{t-s})\right)$
$=(y-x-1)(t-s+1)N^{t-s}+o(N^{t-s}),$
@@
$\frac{(y-s)_{t-s}}{y}-\frac{(x+1-s)_{t-s}}{x+1}$ $=(y-x-1)\psi_{2}^{\prime}(z_{2})$
$=(y-x-1)\left((t-s-1)z_{2}^{t-s-2}+o(z^{t-s-2})\right)$
$=(y-x-1)(t-s-1)N^{t-s-2}+o(N^{t-s-2}).$
Consequently, we have
$g(x,y)$ $=s(y-x-1)(t-s+1)N^{t-s}+t(x+1)y(y-x-1)(t-s-1)N^{t-s-2}+o(N^{t-s})$
$=\left(t+s-(t-s)^{2}\right)(y-x-1)N^{t-s}+o(N^{t-s}),$
which proves Claim 5.2. ∎
It follows from Claim 5.2 that
$\tbinom{x+1}{t}\tbinom{y-1}{s}+\tbinom{x+1}{s}\tbinom{y-1}{t}-\tbinom{x}{t}\tbinom{y}{s}-\tbinom{x}{s}\tbinom{y}{t}$ $=\frac{(x)_{s-1}(y-1)_{s-1}}{t!\,s!}\cdot g(x,y)$
$=\frac{t+s-(t-s)^{2}}{s!t!}\bigl(y-x-1\bigr)N^{t+s-2}+o(N^{t+s-2}),$
which completes the proof of Lemma 5.1. ∎
Lemma 5.3.
Let $F=K_{a_{1},\ldots,a_{r}}$ be an almost balanced complete $r$-partite graph with $a_{1}\geq\cdots\geq a_{r}$, and let $\ell\coloneqq a_{1}+\cdots+a_{r}$. Let $m\coloneqq m_{r,\ell}$ be the constant given by Lemma 3.2. Then
$\binom{\ell}{2}>m\sum_{k\in[r]}\binom{a_{k}}{2}$.
@@
Suppose that $a_{r}=1$. Then it follows from Fact 2.7 (i) that $a_{1}=\cdots=a_{\ell-r}=2$ and $a_{\ell-r+1}=\cdots=a_{r}=1$. Combining this with $m<\frac{\ell(\ell-1)}{2(\ell-r)}$ (by Lemma 4.2 (ii)), we obtain
$m\sum_{k\in[r]}\binom{a_{k}}{2}=m(\ell-r)<\frac{\ell(\ell-1)}{2(\ell-r)}(\ell-r)=\binom{\ell}{2}$.
Now suppose that $a_{r}\geq 2$ (consequently, $\ell\geq 2r$). Observe that
$r\sum_{k\in[r]}\binom{a_{k}}{2}=\frac{1}{2}\Big(r\sum_{k\in[r]}a_{k}^{2}-r\sum_{k\in[r]}a_{k}\Big)$ $=\frac{1}{2}\Big((a_{1}+\cdots+a_{r})^{2}+\sum_{\{i,j\}\in\binom{[r]}{2}}(a_{i}-a_{j})^{2}-r\ell\Big)$
$=\frac{1}{2}\Big(\ell^{2}+\sum_{\{i,j\}\in\binom{[r]}{2}}(a_{i}-a_{j})^{2}-r\ell\Big).$ (34)
For every $\{i,j\}\subseteq[r]$, it follows from the almost balanced assumption that
$\binom{a_{i}-a_{j}}{2}\leq\binom{a_{1}-a_{r}}{2}\leq a_{r}-1\leq a_{j}-1,$
which implies that
$(a_{i}-a_{j})^{2}\leq a_{i}+a_{j}-2.$ (35)
First, we consider the case $r\leq 3$. Since $\ell\geq 2r\geq\frac{\ln(r+1)}{\ln(1+1/r)}$ for $r\in\{2,3\}$, the integer $m_{r,\ell}$ given by Lemma 3.2 satisfies that $m_{r,\ell}=r$. Combining this with (5) and (35), we obtain
$m\sum_{k\in[r]}\binom{a_{k}}{2}=r\sum_{k\in[r]}\binom{a_{k}}{2}$ $=\frac{1}{2}\Big(\ell^{2}+\sum_{\{i,j\}\in\binom{[r]}{2}}(a_{i}-a_{j})^{2}-r\ell\Big)\leq\frac{1}{2}\Bigl(\ell^{2}+\sum_{\{i,j\}\in\binom{[r]}{2}}(a_{i}+a_{j}-2)-r\ell\Bigr)$
$=\frac{1}{2}\bigl(\ell^{2}+(r-1)\ell-r(r-1)-r\ell\bigr)=\frac{1}{2}\bigl(\ell^{2}-\ell-r(r-1)\bigr)<\binom{\ell}{2}$,
as desired. So it remains to prove the case when $r\geq 4$.
Claim 5.4.
We have $\sum_{\{i,j\}\in\binom{[r]}{2}}(a_{i}-a_{j})^{2}\leq(r-1)\ell-r^{2}$.
Proof.
Observe that
$\sum_{\{i,j\}\in\binom{[r]}{2}}(a_{i}-a_{j})^{2}$ $=\sum_{\{i,j\}\in\binom{[2,r-1]}{2}}(a_{i}-a_{j})^{2}+(a_{1}-a_{r})^{2}+\sum_{i\in[2,r-1]}\big((a_{1}-a_{i})^{2}+(a_{i}-a_{r})^{2}\big)$
$\leq\sum_{\{i,j\}\in\binom{[2,r-1]}{2}}(a_{i}-a_{j})^{2}+(a_{1}-a_{r})^{2}+\sum_{i\in[2,r-1]}(a_{1}-a_{r})^{2}$
$=\sum_{\{i,j\}\in\binom{[2,r-1]}{2}}(a_{i}-a_{j})^{2}+(r-1)(a_{1}-a_{r})^{2},$
where the inequality follows from the fact that, for all $x\geq y\geq z\geq 0$,
$(x-y)^{2}+(y-z)^{2}=\big((x-y)+(x-z)\big)^{2}-2(x-y)(y-z)\leq(x-z)^{2}.$
Repeating this argument for $\sum_{{i,j}\in\binom{[2,r-1]}{2}}(a_{i}-a_{j})^{2}$, and continuing similarly, then applying (35), we obtain
$\sum_{\{i,j\}\in\binom{[r]}{2}}(a_{i}-a_{j})^{2}$ $\leq(r-1)(a_{1}-a_{r})^{2}+(r-3)(a_{2}-a_{r-1})^{2}+\cdots+\big(r-2\lfloor\tfrac{r}{2}\rfloor+1\big)\big(a_{\lfloor r/2\rfloor}-a_{r+1-\lfloor r/2\rfloor}\big)^{2}$
$=\sum_{i=1}^{\lfloor r/2\rfloor}(r-2i+1)(a_{i}-a_{r+1-i})^{2}\leq\sum_{i=1}^{\lfloor r/2\rfloor}(r-2i+1)(a_{i}+a_{r+1-i}-2).$
So it suffices to show that
$\sum_{i=1}^{\lfloor r/2\rfloor}(r-2i+1)(a_{i}+a_{r+1-i}-2)\leq(r-1)\ell-r^{2}.$
Suppose that $r=2k$ for some $k\geq 2$. Then
$\sum_{i=1}^{\lfloor r/2\rfloor}(r-2i+1)(a_{i}+a_{r+1-i}-2)$ $=\sum_{i\in[k]}(r-1)(a_{i}+a_{r+1-i}-2)-\sum_{i\in[k]}(2i-2)(a_{i}+a_{r+1-i}-2)$
$\leq(r-1)(\ell-r)-\sum_{i\in[k]}(2i-2)(2+2-2)$
$=(r-1)(\ell-r)-\frac{r(r-2)}{2}=(r-1)\ell-r^{2}-\frac{r(r-4)}{2}\leq(r-1)\ell-r^{2},$
as desired.
Now suppose that r=2k1r=2k-1 for some $k\geq 3$. Then
$\sum_{i=1}^{\lfloor r/2\rfloor}(r-2i+1)(a_{i}+a_{r+1-i}-2)$
$\quad=\sum_{i\in[k-1]}(r-1)(a_{i}+a_{r+1-i}-2)-\sum_{i\in[k-1]}(2i-2)(a_{i}+a_{r+1-i}-2)$
$\quad=\sum_{i=1}^{r}(r-1)(a_{i}-1)-(r-1)(a_{k}-1)-\sum_{i\in[k-1]}(2i-2)(a_{i}+a_{r+1-i}-2)$
$\quad\leq(r-1)(\ell-r)-(r-1)(2-1)-\sum_{i\in[k-1]}(2i-2)(2+2-2)$
$\quad=(r-1)(\ell-r)-(r-1)-\frac{1}{2}\left(r^{2}-4r+3\right)=(r-1)\ell-r^{2}-\frac{r(r-4)+1}{2}\leq(r-1)\ell-r^{2},$
also as desired. This completes the proof of Claim 5.4. ∎
It follows from (5), Claim 5.4, and the inequality $m<\frac{r\ell^{2}}{\ell^{2}-r^{2}}$ (by Lemma 4.2 (ii)) that
$r\sum_{k\in[r]}\tbinom{a_{k}}{2}=\frac{m}{r}\cdot r\sum_{k\in[r]}\tbinom{a_{k}}{2}$ $=\frac{m}{r}\cdot\frac{1}{2}\Big(\ell^{2}+\sum_{\{i,j\}\in\binom{[r]}{2}}(a_{i}-a_{j})^{2}-r\ell\Big)$
$\leq\frac{\ell^{2}}{\ell^{2}-r^{2}}\Big(\ell^{2}+(r-1)\ell-r^{2}-r\ell\Big)=\binom{\ell}{2}-\frac{r^{2}\ell}{2(\ell^{2}-r^{2})}<\binom{\ell}{2}$.
This completes the proof of Lemma 5.3. ∎
@@
Let $F=K_{a_{1},\ldots,a_{r}}$ be an almost balanced complete $r$-partite graph on $\ell$ vertices, that is, $\ell=a_{1}+\cdots+a_{r}$. Let $m=m_{r,\ell}$ be the integer given by Lemma 3.2. Let $n$ be a sufficiently large integer, and let $N\coloneqq n/m$. Let $G$ be an $n$-vertex graph satisfying $I(F,G)=I(F,n)$.
It follows from Theorems 3.1 and 1.5 that $G$ is complete $m$-partite with each part of size $(1+o(1))N$. Write the part sizes of $G$ as $n_{1}\geq n_{2}\geq\dots\geq n_{m}$, noting that $\sum_{i=1}^{m}n_{i}=n$ and $|n_{i}-N|=o(n)$ for every $i\in[m]$. Let $y\coloneqq n_{1}$ and $x\coloneqq n_{m}$. Suppose to the contrary that $G$ is not a Turán graph, that is, $y\geq x+2$.
For convenience, for every graph $H$, let $\hat{I}(F,H)\coloneqq\mathrm{sym}(a_{1},\dots,a_{r})\cdot I(F,H)$.
@@
The complete $m$-partite graph $G_{\ast}\cong K_{n_{1}-1,n_{2},\dots,n_{m}+1}$ satisfies
$\hat{I}(F,G_{\ast})-\hat{I}(F,G)>0.$
@@
Recall that $m=2$ when $r=2$ by Lemma 3.2. For each $i\in[r]$, let
$f_{i}(x,y)$ $\coloneqq\tbinom{x+1}{a_{i}}+\tbinom{y-1}{a_{i}}-\tbinom{x}{a_{i}}-\tbinom{y}{a_{i}}$,
@@
$A_{i}$ $\coloneqq\begin{cases}0,&\quad\text{if}\quad r=2,\\[3.00003pt] \sum_{(s_{k})_{k\in[r]\setminus\{i\}}}\prod_{k\in[r]\setminus\{i\}}\tbinom{n_{s_{k}}}{a_{k}},&\quad\text{if}\quad r\geq 3,\end{cases}\quad=\left(\tfrac{(m-2)_{r-1}\cdot a_{i}!}{a_{1}!\cdots a_{r}!}+o(1)\right)N^{\ell-a_{i}}.$
where the tuple $(s_{k})_{k\in[r]\setminus\{i\}}$ ranges over $\big([2,m-1]\big)_{r-1}$ in the summation.
For each $\{i,j\}\subseteq[r]$, let
$g_{i,j}(x,y)$ $\coloneqq\tbinom{x+1}{a_{i}}\tbinom{y-1}{a_{j}}+\tbinom{x+1}{a_{j}}\tbinom{y-1}{a_{i}}-\tbinom{x}{a_{i}}\tbinom{y}{a_{j}}-\tbinom{x}{a_{j}}\tbinom{y}{a_{i}}$,
@@
$B_{i,j}$ $\coloneqq\begin{cases}1,&\quad\text{if}\quad r=2,\\[3.00003pt] \sum_{(s_{k})_{k\in[r]\setminus\{i,j\}}}\prod_{k\in[r]\setminus\{i,j\}}\tbinom{n_{s_{k}}}{a_{k}},&\quad\text{if}\quad r\geq 3,\end{cases}\quad=\left(\tfrac{(m-2)_{r-2}\cdot a_{i}!a_{j}!}{a_{1}!\cdots a_{r}!}+o(1)\right)N^{\ell-a_{i}-a_{j}},$
where the tuple $(s_{k})_{k\in[r]\setminus\{i,j\}}$ ranges over $\big([2,m-1]\big)_{r-2}$ in the summation.
Since $\mathrm{sc}(G)=0$, applying (4) and simplifying, we obtain
$\hat{I}(F,G_{\ast})-\hat{I}(F,G)$ $=\sum_{i\in[r]}A_{i}\cdot f_{i}(x,y)+\sum_{\{i,j\}\in\binom{[r]}{2}}B_{i,j}\cdot g_{i,j}(x,y).$ (36)
It follows from (32) that
$A_{i}\cdot f_{i}(x,y)$ $=-\frac{a_{i}(a_{i}-1)}{a_{i}!}(y-x-1+o(1))N^{a_{i}-2}\cdot\left(\frac{(m-2)_{r-1}\cdot a_{i}!}{a_{1}!\cdots a_{r}!}+o(1)\right)N^{\ell-a_{i}}$
$=a_{i}(a_{i}-1)(y-x-1)\frac{(m-2)_{r-1}}{a_{1}!\cdots a_{r}!}N^{\ell-2}+o(N^{\ell-2}).$ (37)
It follows from (33) that
$B_{i,j}\cdot g_{i,j}(x,y)$ $=\frac{a_{i}+a_{j}-(a_{i}-a_{j})^{2}}{a_{i}!a_{j}!}\big(y-x-1+o(1)\big)N^{a_{i}+a_{j}-2}\cdot\left(\frac{(m-2)_{r-2}\cdot a_{i}!a_{j}!}{a_{1}!\cdots a_{r}!}+o(1)\right)N^{\ell-a_{i}-a_{j}}$
$=\left(a_{i}+a_{j}-(a_{i}-a_{j})^{2}\right)(y-x-1)\frac{(m-2)_{r-2}}{a_{1}!\cdots a_{r}!}N^{\ell-2}+o(N^{\ell-2}).$ (38)
Let
$\Psi\coloneqq-\sum_{i\in[r]}(m-r)a_{i}(a_{i}-1)+\sum_{\{i,j\}\in\binom{[r]}{2}}\left(a_{i}+a_{j}-(a_{i}-a_{j})^{2}\right).$
Then
$\Psi$ $=-\sum_{i\in[r]}(m-r)(a_{i}^{2}-a_{i})+\sum_{\{i,j\}\in\binom{[r]}{2}}\left(a_{i}+a_{j}-a_{i}^{2}-a_{j}^{2}+2a_{i}a_{j}\right)$
$=\sum_{i\in[r]}\left(-(m-r)(a_{i}^{2}-a_{i})-(r-1)(a_{i}^{2}-a_{i})\right)+\sum_{i\in[r]}a_{i}\sum_{j\in[r]\setminus\{i\}}a_{j}$
$=-\sum_{i\in[r]}(m-1)(a_{i}^{2}-a_{i})+\sum_{i\in[r]}a_{i}(\ell-a_{i})$
$=-\sum_{i\in[r]}m(a_{i}^{2}-a_{i})+\sum_{i\in[r]}(\ell-1)a_{i}=2\Big(\tbinom{\ell}{2}-m\sum_{i\in[r]}\tbinom{a_{i}}{2}\Big)$.
Combining this with (36), (5), and (5), we obtain
$\hat{I}(F,G_{\ast})-\hat{I}(F,G)$ $=\Psi\cdot(y-x-1)\frac{(m-2)_{r-2}}{a_{1}!\cdots a_{r}!}N^{\ell-2}+o(N^{\ell-2})$
$=2\Big(\tbinom{\ell}{2}-m\sum_{i\in[r]}\tbinom{a_{i}}{2}\Big)\cdot(y-x-1)\frac{(m-2)_{r-2}}{a_{1}!\cdots a_{r}!}N^{\ell-2}+o(N^{\ell-2}).$
By Lemma 5.3, the factor $\binom{\ell}{2}-m\sum_{i}\binom{a_{i}}{2}$ is positive and hence at least one. It follows that $\hat{I}(F,G_{\ast})-\hat{I}(F,G)>0$, which establishes Claim 5.5. ∎
It follows from Claim 5.5 that $G$ is not extremal, contradicting the assumption that $I(F,G)=I(F,n)$. This completes the proof of Theorem 1.1. ∎
@@
Let $r,t\geq 2$ be integers. The following holds for every $n\geq rt$. Suppose that $G$ is an $n$-vertex complete multipartite graph satisfying $I(K_{r}(t),G)=I(K_{r}(t),n)$. Then $G\cong T_{k}(n)$ for some integer $k\geq r$. In particular, for every integer $n\geq rt$, there exists an integer $k$ such that $I(K_{r}(t),T_{k}(n))=I(K_{r}(t),n)$.
@@
Fix integers $r\geq 2$ and $t\geq 2$, and let $F=K_{r}(t)$. Fix $n\geq rt$. By Lemma 2.3, there exists a complete multipartite graph $G$ such that $I(F,G)=I(F,n)$. Write $G=K_{a_{1},\ldots,a_{k}}$ for some integer $k$, and assume that $a_{1}\geq\cdots\geq a_{k}$. Clearly, we must have $k\geq r$ and $a_{k}\geq t$, since otherwise $I(F,G)=0$.
Suppose to the contrary that $a_{1}\geq a_{k}+2$. Let $G^{\prime}$ be the complete $k$-partite graph with part sizes $a_{1}-1,a_{2},\cdots,a_{k-1},a_{k}+1$, and let $G^{\prime\prime}$ be the complete (k1)(k-1)-partite graph with part sizes $a_{1}+a_{k},a_{2},\cdots,a_{k-1}$.
@@
We have $I(F,G)<\max\{I(F,G^{\prime}),~I(F,G^{\prime\prime})\}$.